Given that 2π‘₯ βˆ’ 5 is a factor of polynomial functions 𝑓(π‘₯) = π‘˜π‘₯^2 βˆ’ 𝑏π‘₯ + 3π‘˜, determine the ratio k : b in the simplest form.


1
Expert's answer
2021-11-24T10:15:27-0500

If "2\ud835\udc65 \u2212 5" is a factor of polynomial functions "\ud835\udc53(\ud835\udc65) = \ud835\udc58\ud835\udc65^2 \u2212 \ud835\udc4f\ud835\udc65 + 3\ud835\udc58," then


"𝑓(\dfrac{5}{2}) = 𝑘(\dfrac{5}{2})^2 − 𝑏(\dfrac{5}{2}) + 3𝑘=0"

"25k-10b+12k=0"

"37k=10b"

"k:b=10:37"